OpenStudy (anonymous):

What is the missing number? Please explain and don't give an answer. 4, 13, 40, __, 364, 1093

OpenStudy (anonymous):

How you figured it out is just fine! :)

OpenStudy (anonymous):

3n + 1 40 = 3*13 + 1 so next number is 3* -- you can do this one

OpenStudy (campbell_st):

look at the differences 13 - 4 = 9 = 9^2 40 - 13 = 27 = 9^3 1093 = 464 = 729 = 3^6

OpenStudy (campbell_st):

so add 3^n+1 to the previous term. n the the term number

OpenStudy (anonymous):

What am I adding it to?

OpenStudy (campbell_st):

4, 4+3^2, (4 + 3^2)+3^3, .(4+3^2+3^3)+3^4 etc

OpenStudy (anonymous):

OK. 121?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Thanks!